Solving\\(x^{2})=3\\
1
Expert's answer
2019-11-13T09:29:10-0500

Solving Quadratic Equation


We need to find the roots of the Quadratic equation "x^2 = 3"


Solution:


The given Quadratic equation is

"x^2 = 3"

Apply the square root on both sides


"\sqrt {x^2} = \pm \sqrt 3"

"x = \pm \sqrt 3"

Answer:

Solutions of the given equation is "x = \pm \sqrt 3"